OpenStudy (anonymous):

PLEASE HELP ME!! FAN AND MEDAL City A and City B had two different temperatures on a particular day. On that day, four times the temperature of City A was 8° C more than 3 times the temperature of City B. The temperature of City A minus twice the temperature of City B was –3° C. What was the temperature of City A and City B on that day?

OpenStudy (anonymous):

City A and City B had two different temperatures on a particular day. On that day, four times the temperature of City A was 8° C more than 3 times the temperature of City B. The temperature of City A minus twice the temperature of City B was –3° C. What was the temperature of City A and City B on that day? City A was 5° C, and City B was 4° C. City A was 3° C, and City B was –1° C. City A was 8° C, and City B was –3° C. City A was 5° C, and City B was –5° C.

OpenStudy (anonymous):

@amistre

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

city A was 8 degrees c and city b was -3 degrees c

OpenStudy (anonymous):

@itsbribro @ganeshie8

OpenStudy (amistre64):

we have 2 equations that we can construct with this

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

equation 1 is: A is 8 more than, 3 times B equation 2 is: A - 2B is -3

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

i think the answer is c

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

A = 8 + 3B A-2B = -3 subtract one from the other to get A = 8 + 3B -A+2B = +3 --------------- 2B = 11 + 3B hmm, im getting B = -11 gotta chk to see that im reading it right to start with then

OpenStudy (amistre64):

4ties A is 8 more than .... so 4A = 8 + 3B A-2B = -3 4A = 8 + 3B -4A+8B = 12 ------------------ 8B = 20 + 3B thats better

OpenStudy (help_people):

@amistre64 srry, but just making sure this is c right?

OpenStudy (amistre64):

8B = 20 + 3B , solve for B

OpenStudy (help_people):

B=-3 is what I found am I correct?

OpenStudy (amistre64):

lets check and see 8B = 20 + 3B 8(-3) = 20 + 3(-3) -24 = 20 - 9 -24 = 11 nope, not going to work out is it

OpenStudy (help_people):

no

OpenStudy (help_people):

B=-1 maybe

OpenStudy (fatimab2000):

no its a